# Spectral Moment Formulae for $\hbox{GL}(3)\times\hbox{GL}(2)$ L-functions II: The Eisenstein Case

### v2

@@ 7.1. Agreement with the unitary CFKRS
Notice that the line of integration is shifted by a small quantity $\gamma$. We put $s:=(1+\delta)/2$, and apply the functional equation of the $\zeta$-function to the factor $\zeta(s_{0}+\nu_{3})$, the moment (7.1) now takes the form:
@@
$\hskip 20.0pt\ \cdot\ \frac{\Gamma_{\mathbb{R}}(1/2-\gamma-\nu_{3}-it)}{\Gamma_{\mathbb{R}}(1/2+\gamma+\nu_{3}+it)}\ \left(\mathcal{F}_{\nu}H\right)\Big(\frac{1}{2}+\gamma+it,\ \frac{1+\delta}{2}\Big)\,dt.$ (7.2)
Following Section 5.2, we introduce the linear re-parametrizations:
$\begin{cases}\alpha_{1}\ =&\gamma\ +\ \nu_{1}\\ \alpha_{2}\ =&\gamma\ +\ \nu_{2}\\ \beta_{1}\ =&-\gamma\ -\ \nu_{3}\\ \beta_{2}\ =&\delta\ -\ \gamma,\end{cases}$ (7.3)
@@
$\int_{\mathbb{R}}\,\zeta\Big(\frac{1}{2}+\alpha_{1}+it\Big)$ $\zeta\Big(\frac{1}{2}+\alpha_{2}+it\Big)\zeta\Big(\frac{1}{2}+\beta_{1}-it\Big)\zeta\Big(\frac{1}{2}+\beta_{2}-it\Big)\eta_{\alpha_{1},\alpha_{2};\beta_{1},\beta_{2}}(t)\,dt$ (7.5)
with the weight function given by
@@
We write $\mathcal{R}_{\alpha}(s;H)=\mathcal{R}_{\nu}(s;H)$ (see (1.16)) and $\mathcal{D}_{\alpha}(s;H)=\mathcal{D}_{\nu}(s;H)$ (see (1.15)) using the re-labeling (7.1). We are ready to compare them with Conjecture (5.2). Firstly, we have
$\frac{1}{2}\,\mathcal{R}_{\nu}(s;H)\ =\$$H(1-s+\nu_{1})\frac{\Lambda\left(1+\nu_{1}-\nu_{2}\right)\Lambda(2s-1-\nu_{1}-\nu_{2})\Lambda\left(1+\nu_{1}-\nu_{3}\right)\Lambda(2s-1-\nu_{1}-\nu_{3})}{\Lambda\left(3-2s+2\nu_{1}\right)}$
$\ +\ H(1-s+\nu_{2})\frac{\ \Lambda\left(1+\nu_{2}-\nu_{1}\right)\Lambda(2s-1-\nu_{2}-\nu_{1})\Lambda\left(1+\nu_{2}-\nu_{3}\right)\Lambda(2s-1-\nu_{2}-\nu_{3})}{\Lambda\left(3-2s+2\nu_{2}\right)}$
$\hskip 10.0pt\ +\ H(1-s+\nu_{3})\frac{\Lambda\left(1+\nu_{3}-\nu_{1}\right)\Lambda(2s-1-\nu_{3}-\nu_{1})\Lambda\left(1+\nu_{3}-\nu_{2}\right)\Lambda(2s-1-\nu_{3}-\nu_{2})}{\Lambda\left(3-2s+2\nu_{3}\right)}.$
Then by (7.4) and repeated use of the functional equation of $\Lambda(s)$, it follows that
$\frac{1}{2}\,\mathcal{R}_{\nu}(s;H)\ =\$$H\Big(\frac{1+\alpha_{1}-\alpha_{2}+\beta_{1}-\beta_{2}}{2}\Big)\frac{\Lambda(1+\alpha_{1}-\alpha_{2})\Lambda(1+\beta_{1}-\beta_{2})\Lambda(1+\alpha_{1}+\beta_{1})\Lambda(1-\alpha_{2}-\beta_{2})}{\Lambda(2+\alpha_{1}-\alpha_{2}+\beta_{1}-\beta_{2})}$
$\ +\ H\Big(\frac{1-\alpha_{1}+\alpha_{2}+\beta_{1}-\beta_{2}}{2}\Big)\frac{\Lambda(1-\alpha_{1}+\alpha_{2})\Lambda\left(1+\beta_{1}-\beta_{2}\right)\Lambda(1+\alpha_{2}+\beta_{1})\Lambda(1-\alpha_{1}-\beta_{2})}{\Lambda(2-\alpha_{1}+\alpha_{2}+\beta_{1}-\beta_{2})}$
$\hskip 10.0pt\ +\ H\Big(\frac{1-\alpha_{1}-\alpha_{2}-\beta_{1}-\beta_{2}}{2}\Big)\frac{\Lambda(1-\alpha_{1}-\beta_{1})\Lambda(1-\alpha_{2}-\beta_{2})\Lambda(1-\alpha_{2}-\beta_{1})\Lambda(1-\alpha_{1}-\beta_{2})}{\Lambda(2-\alpha_{1}-\alpha_{2}-\beta_{1}-\beta_{2})}.$
Secondly, recall the Weyl actions of $w_{3},w_{5},w_{\ell}$ on the parameters $(\nu_{1},\nu_{2},\nu_{3})$ in (1.4). Then we have
$\frac{1}{2}\ \mathcal{D}_{\nu}(s;H)\ =\$$H(s-\nu_{1})\ \frac{\Lambda\left(1-\nu_{1}+\nu_{3}\right)\Lambda\left(1-\nu_{1}+\nu_{2}\right)\Lambda\left(2s-\nu_{1}-\nu_{3}\right)\Lambda\left(2s-\nu_{1}-\nu_{2}\right)}{\Lambda(1+2(s-\nu_{1}))}$
$\ +\ H(s-\nu_{2})\ \frac{\Lambda\left(1-\nu_{2}+\nu_{3}\right)\Lambda\left(1-\nu_{2}+\nu_{1}\right)\Lambda\left(2s-\nu_{2}-\nu_{3}\right)\Lambda\left(2s-\nu_{2}-\nu_{1}\right)}{\Lambda(1+2(s-\nu_{2}))}$
$\hskip 10.0pt\ +\ H(s-\nu_{3})\ \frac{\Lambda\left(1-\nu_{3}+\nu_{2}\right)\Lambda\left(1-\nu_{3}+\nu_{1}\right)\Lambda\left(2s-\nu_{3}-\nu_{2}\right)\Lambda\left(2s-\nu_{3}-\nu_{1}\right)}{\Lambda(1+2(s-\nu_{3}))}.$
With (7.4), we obtain the following:
$\frac{1}{2}\ \mathcal{D}_{\nu}(s;H)\ =\$$H\Big(\frac{1-\alpha_{1}+\alpha_{2}-\beta_{1}+\beta_{2}}{2}\Big)\frac{\Lambda(1-\alpha_{1}-\beta_{1})\Lambda(1-\alpha_{1}+\alpha_{2})\Lambda(1+\alpha_{2}+\beta_{2})\Lambda(1-\beta_{1}+\beta_{2})}{\Lambda(2-\alpha_{1}+\alpha_{2}-\beta_{1}+\beta_{2})}$
$\ +\ H\Big(\frac{1+\alpha_{1}-\alpha_{2}-\beta_{1}+\beta_{2}}{2}\Big)\frac{\Lambda(1-\alpha_{2}-\beta_{1})\Lambda(1+\alpha_{1}-\alpha_{2})\Lambda(1+\alpha_{1}+\beta_{2})\Lambda(1-\beta_{1}+\beta_{2})}{\Lambda(2+\alpha_{1}-\alpha_{2}-\beta_{1}+\beta_{2})}$
$\hskip 10.0pt\ +\ H\Big(\frac{1+\alpha_{1}+\alpha_{2}+\beta_{1}+\beta_{2}}{2}\Big)\frac{\Lambda(1+\alpha_{2}+\beta_{1})\Lambda(1+\alpha_{1}+\beta_{1})\Lambda(1+\alpha_{1}+\beta_{2})\Lambda(1+\alpha_{2}+\beta_{2})}{\Lambda(2+\alpha_{1}+\alpha_{2}+\beta_{1}+\beta_{2})}.$
Now, we readily observe that the quotients and products of $\zeta$’s in $\mathcal{R}_{\nu}(s;H)+\mathcal{D}_{\nu}(s;H)$ match up exactly with those in Conjecture (5.2)! Moreover, the cubic moment dual to (7.5) takes the form:
$\sum_{j}\ H(\mu_{j})\,\frac{\Lambda\big(\frac{1-\alpha_{1}+\alpha_{2}-\beta_{1}+\beta_{2}}{2},\,\phi_{j}\big)\Lambda\big(\frac{1-\alpha_{2}+\alpha_{1}-\beta_{1}+\beta_{2}}{2},\,\phi_{j}\big)\Lambda\big(\frac{1+\alpha_{1}+\alpha_{2}+\beta_{1}+\beta_{2}}{2},\,\phi_{j}\big)}{\langle\,\phi_{j},\phi_{j}\,\rangle}$,
@@ 7.2. Comments
In the literature, moments of $L$-functions are more commonly approached with the approximate functional equations. A clear advantage is that one can take $s=1/2$ right from the start. For example, tracing the arguments of [Pe15] for the cubic moment of $\hbox{GL}(2)$ $L$-functions (readily adaptable to the Maass case and closely related to [Y17]), the main terms for the fourth moment of the $\hbox{GL}(1)$ $L$-functions do not appear: most likely, they are small and absorbed into the error term of the approximate Motohashi-type formula obtained in [Pe15]. In other words, the “approximate” treatment precludes the possibility of spectral inversion by changing the test vectors (see the end of Section 1.5). This is not satisfactory, as one should be able to pass between the two different-looking moments in a Motohashi-type formula for distinct applications, much as with the celebrated Kuznetsov formulae.
In Section 1.4, we compare the sources of the main terms in the fourth moment with earlier methods. For the cubic moment, previous works of [Iv02] and [Fr20] rely on the shifted divisor sums to extract the main terms. The work [Fr20] (in the weight aspect) is closer in spirit to [CFK+05], which begins with an exact identity for the twisted second moment of $\hbox{GL}(2)$ $L$-functions, obtained quite non-trivially from the Petersson formula. The dual side of the moment identity comprises three pieces of shifted divisor sums. In total, $12$ different residues arise from these sums. As in [HY10, Y11], delicate combinations of terms are required, and [Fr20] carries this out. Interestingly, some combinations contribute to main terms, some to error terms, and some to a mixture of both! This appears to be the primary reason [Fr20] arrives at an approximate cubic moment identity, despite starting with an exact moment identity. The approaches of [Mo93, BM05, Ne20+, Wu22, BFW21+] and the present work do not encounter this issue. This phenomenon deserves further investigation.

### v3

@@ 7.1. Agreement with the unitary CFKRS
Notice that the line of integration is shifted by a small quantity $\gamma$. We put
$s\,:=\,\frac{1+\delta}{2},$ (7.2)
and apply the functional equation (2.3) of the $\zeta$-function to the factor $\zeta(s_{0}+\nu_{3})$, the moment (7.1) takes the form:
@@
$\hskip 20.0pt\ \cdot\ \frac{\Gamma_{\mathbb{R}}(1/2-\gamma-\nu_{3}-it)}{\Gamma_{\mathbb{R}}(1/2+\gamma+\nu_{3}+it)}\ \left(\mathcal{F}_{\nu}H\right)\Big(\frac{1}{2}+\gamma+it,\ \frac{1+\delta}{2}\Big)\,dt.$
Following the setting of Section 5.2, we introduce the linear re-parametrizations:
$\begin{cases}\alpha_{1}\ =&\gamma\ +\ \nu_{1}\\ \alpha_{2}\ =&\gamma\ +\ \nu_{2}\\ \beta_{1}\ =&-\gamma\ -\ \nu_{3}\\ \beta_{2}\ =&\delta\ -\ \gamma,\end{cases}$
@@
$\int_{\mathbb{R}}\,\zeta\Big(\frac{1}{2}+\alpha_{1}+it\Big)$ $\zeta\Big(\frac{1}{2}+\alpha_{2}+it\Big)\zeta\Big(\frac{1}{2}+\beta_{1}-it\Big)\zeta\Big(\frac{1}{2}+\beta_{2}-it\Big)\eta_{\alpha_{1},\alpha_{2};\beta_{1},\beta_{2}}(t)\,dt,$ (7.4)
where the weight function is given by
@@
We write $\mathcal{R}_{\pm\alpha}(s;H)=\mathcal{R}_{\pm\nu}(s;H)$ (see (1.13)). Using the re-labeling (7.3) and repeated application of the functional equation of $\Lambda(s)$, it follows that $(1/2)\,\mathcal{R}_{\nu}(1-s;H)+(1/2)\,\mathcal{R}_{-\nu}(s;H)$ is given by
$H\Big(\frac{1+\alpha_{1}-\alpha_{2}+\beta_{1}-\beta_{2}}{2}\Big)\frac{\Lambda(1+\alpha_{1}-\alpha_{2})\Lambda(1+\beta_{1}-\beta_{2})\Lambda(1+\alpha_{1}+\beta_{1})\Lambda(1-\alpha_{2}-\beta_{2})}{\Lambda(2+\alpha_{1}-\alpha_{2}+\beta_{1}-\beta_{2})}$
[…]
$+$ $H\Big(\frac{1-\alpha_{1}+\alpha_{2}-\beta_{1}+\beta_{2}}{2}\Big)\frac{\Lambda(1-\alpha_{1}-\beta_{1})\Lambda(1-\alpha_{1}+\alpha_{2})\Lambda(1+\alpha_{2}+\beta_{2})\Lambda(1-\beta_{1}+\beta_{2})}{\Lambda(2-\alpha_{1}+\alpha_{2}-\beta_{1}+\beta_{2})}$
[…]
We now readily observe that the quotients and products of $\zeta$’s above match exactly with those in the CFKRS Conjecture (5.2)! Moreover, the cubic moment dual to (7.4) takes the form:
$\sum_{j}\ \frac{H(\mu_{j})}{\langle\,\phi_{j},\phi_{j}\,\rangle}\,\Lambda\big(\frac{1-\alpha_{1}+\alpha_{2}-\beta_{1}+\beta_{2}}{2},\,\phi_{j}\big)\Lambda\big(\frac{1-\alpha_{2}+\alpha_{1}-\beta_{1}+\beta_{2}}{2},\,\phi_{j}\big)\Lambda\big(\frac{1+\alpha_{1}+\alpha_{2}+\beta_{1}+\beta_{2}}{2},\,\phi_{j}\big)$
@@ 7.2. Comments
In the literature, moments of $L$-functions are more commonly approached with the approximate functional equations. A clear advantage is that one can take $s=1/2$ right from the start. For example, tracing the arguments of [64] for the cubic moment of $\hbox{GL}(2)$ $L$-functions (readily adaptable to the Maass case and closely related to [74]), the main terms for the fourth moment of the $\hbox{GL}(1)$ $L$-functions do not appear: most likely, they are small and absorbed into the error term of the approximate Motohashi-type formula obtained in [64]. In other words, the “approximate” treatment precludes the possibility of spectral inversion by changing the test vectors (see the end of Section 1.5). This is not satisfactory, as one should be able to pass between the two different-looking moments in a Motohashi-type formula for distinct applications, much as with the celebrated Kuznetsov formulae.
In Section 1.4, we compare the sources of the main terms in the fourth moment with earlier methods. For the cubic moment, previous works of [43] and [30] rely on the shifted divisor sums to extract the main terms. The work [30] (in the weight aspect) is closer in spirit to [22], which begins with an exact identity for the twisted second moment of $\hbox{GL}(2)$ $L$-functions, obtained quite non-trivially from the Petersson formula. The dual side of the moment identity comprises three pieces of shifted divisor sums. In total, $12$ different residues arise from these sums. As in [35, 73], delicate combinations of terms are required, and [30] carries this out. Interestingly, some combinations contribute to main terms, some to error terms, and some to a mixture of both! This appears to be the primary reason [30] arrives at an approximate cubic moment identity, despite starting with an exact moment identity. The approaches of [59, 12, 62, 71, 3] and the present work do not encounter this issue. This phenomenon deserves further investigation.